OpenStudy (anonymous):

What is the slope of the line: y= 3x-5? PLEASE HELP!!!

OpenStudy (anonymous):

3

OpenStudy (anonymous):

for a line given by y=mx+c , m is the slope......hence compare and u'll get m=3

OpenStudy (anonymous):

For a function in the form \(y=mx+b\), the slope is m. In our case \(m=3\).

OpenStudy (anonymous):

THANKS GUYS!! ILL FAN YOU GUYSS

OpenStudy (anonymous):

him1618!!! i answered ur question posted a month ago. check if its rite....its some complex number question. ps:sorry using this thread to chat with him :/

OpenStudy (anonymous):

its also important to know that in Y=mx+b, b is the y-intercept so -5 is your y-intercept.